OpenStudy (anonymous):

You want to make five-letter codes that use the letters A, F, E, R, and M without repeating any letter. What is the probability that a randomly chosen code starts with A? A-0.10 B-0.15 C-0.20 D-0.25

jimthompson5910 (jim_thompson5910):

There are 5 letters to choose from. So you have 5*4*3*2*1 = 120 different ways to arrange them (order matters and no repeats are allowed)

jimthompson5910 (jim_thompson5910):

Now if you make A the first letter, then you have 4 letters to arrange in the 4 remaining slots. So you have 4*3*2*1 = 24 ways to do that.

jimthompson5910 (jim_thompson5910):

Divide the two to get the probability: 24/120 = 1/5 = 0.20